# Connectedness of level sets for non-degenerate integrable systems that extend complexity one torus actions

### v2

@@ 4. Purely elliptic type
Since $p$ has tall local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, Remark 2.3 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$, where $\rho\colon H\to(S^{1})^{h+1}$ is the symplectic slice representation at $p$. Hence, by Lemma 4.4 and the Morse Lemma, after possibly replacing $g$ by $-g$, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, that $g([t,\alpha,z])=|z|^{2}$, and that $p=[1,0,0]$. Let $N$ be the degree of the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$. By Lemma 2.7, the map $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a homeomorphism; moreover, it restricts to a diffeomorphism from $({Y/\!/T})\smallsetminus\{[t,0,0]\}$ to ${\mathbb{C}}\smallsetminus\{0\}$. By Lemma 2.9, there exists a positive real number $C$ such that $g([t,\alpha,z])=C|P(z)|^{\frac{2}{N}}$ for all $[t,\alpha,z]\in\Phi_{Y}^{-1}(0)$. Let $S\colon{\mathbb{C}}\to{\mathbb{C}}$ be the homeomorphism that, in polar coordinates, takes $(r,\theta)$ to $(\sqrt{C}\,r^{\frac{1}{N}},\theta)$; it restricts to a diffeomorphism ${\mathbb{C}}\smallsetminus\{0\}\to{\mathbb{C}}\smallsetminus\{0\}$. Define $\Psi\colon{Y/\!/T}\to{\mathbb{C}}\simeq{\mathbb{R}}^{2}$ by $\Psi([t,\alpha,z])=S(P(z))$. Since
@@
for all $[t,\alpha,z]\in{Y/\!/T}$, the map $\Psi$ satisfies properties (1) – (3). ∎
@@ 5. The semitoric case
The goal of this section is to prove that focus-focus points satisfy the hypotheses of Proposition 3.4 (see Lemma 5.4 below). This allows us to provide a short and direct new proof of Vu Ngoc’s theorem that the level sets of semitoric systems are connected [VN] – see Theorem 5.7. In both proofs, the key arguments follow from Morse theory; the most significant difference is that we construct a new smooth structure on the reduced spaces, while he uses the natural smooth structure on $M$.
@@
Identify $T$ with $S^{1}$. By the local normal form for focus-focus points, we may assume that the integrable system is $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},\varphi\circ f_{Y})$ and $p$ is the origin, where $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},f_{Y}=(\Phi_{Y},g_{Y}))$ is as in Example 5.2 and $\varphi\colon{\mathbb{R}}\times{\mathbb{R}}\to{\mathbb{R}}\times{\mathbb{R}}$ is a diffeomorphism [C, VW]. Since by assumption the first component of $\varphi\circ f_{Y}$ induces an effective $S^{1}$-action on $Y$, and since the level sets of $f_{Y}$ are not compact, up to a translation $\Phi_{Y}$ and the first component of $\varphi\circ f_{Y}$ are equal up to sign (cf. [HSS, Proposition 3.9]). Since the integrable systems $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},f_{Y}=(\Phi_{Y},g_{Y}))$ and $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},(-\Phi_{Y},g_{Y}))$ are isomorphic, the claim follows. ∎
@@
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be an integrable system so that $(M,\omega,\Phi)$ is a 4-dimensional complexity one $T$-space. Assume that $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is focus-focus. Then there exist an open neighborhood $U$ of $[p]\in{M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
@@
We are now almost ready to give our proof of Vu Ngoc’s theorem on the connectedness of the level sets of semitoric systems; however, we also need the following important result, which is proved in [At, GS2, LMTW].
@@
Let $T$ act on a connected symplectic manifold $(M,\omega)$ with moment map $\Phi\colon M\to{\mathfrak{t}}^{*}$. If $\Phi$ is proper then the level set $\Phi^{-1}(\alpha)$ is connected for all $\alpha\in{\mathfrak{t}}^{*}.$
@@
Let $N$ be a connected manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a proper Morse-Bott function. If no critical point of $\overline{g}$ has index or coindex one, then the level set $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$.
@@
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be a semitoric system so that $\Phi$ is proper. Then the level set $f^{-1}(\alpha,c)=\Phi^{-1}(\alpha)\cap g^{-1}(c)$ is connected for all $(\alpha,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$.
@@
Fix $\alpha\in{\mathfrak{t}}^{*}$. Assume without loss of generality that $\alpha=0$. Since $\Phi$ is proper, the reduced space ${M/\!/T}$ is compact and connected by Theorem 5.5.
If the reduced space ${M/\!/T}$ is a single point then obviously $\Phi^{-1}(0)$ consists of a single $T$-orbit. Hence, the level set $f^{-1}(0,c)$ is connected for all $c\in{\mathbb{R}}$.
Otherwise, since the orbit of a short point in $\Phi^{-1}(0)$ is an isolated point in $\Phi^{-1}(0)/T$, every point in $\Phi^{-1}(0)$ is tall. If $p\in\Phi^{-1}(0)$ is a critical point of $g$ modulo $\Phi$, then $p$ is also a singular point of $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$. So, by assumption, $p$ either has purely elliptic type or is focus-focus. By Propositions 3.4 and 4.5, and Lemma 5.4, ${M/\!/T}$ can be given the structure of a smooth closed, connected, oriented surface such that $\overline{g}:{M/\!/T}\to{\mathbb{R}}$ is a Morse function with no critical point of index 1. By Proposition 5.6, the level set $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$. Since $T$ is connected, it follows that $f^{-1}(0,c)$ is also connected. ∎
@@ 6. Non-degenerate singular points
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Let $p\in\Phi^{-1}(0)$ be a tall non-degenerate singular point. By Lemma 3.2 and Proposition 4.5, if $p$ has purely elliptic type, then we can put a smooth structure on ${M/\!/T}$ near $[p]$ so that $\overline{g}$ is locally Morse, and hence $p$ satisfies the hypotheses of Proposition 3.4. The aim of this section is to extend this result to the remaining cases (see Proposition 6.2). In particular, if $p$ either has a focus-focus block, or has a hyperbolic block and disconnected $T$-stabilizer, then $[p]$ is a regular point of $\overline{g}$, while if $p$ has a hyperbolic block and connected $T$-stabilizer, then $[p]$ is a singular point with index one.
We start by recalling some facts about non-degenerate singular points.
@@
We can now state our main proposition.
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. If $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is non-degenerate it has at most one non-elliptic block; moreover,
  • If $p$ has a hyperbolic block and connected $T$-stabilizer, then there exist an open neighborhood $U$ of $[p]\in{M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
@@
      $\overline{g}\circ\Psi^{-1}(x,y)=x^{2}-y^{2}$ for all $(x,y)\in\Psi(U)$.
  • If $p$ either has a focus-focus block, or has a hyperbolic block and disconnected $T$-stabilizer, then there exist an open neighborhood $U$ of $[p]\in{M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:

### v3

@@ 4. Purely elliptic type
Since $p$ has tall local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, Remark 2.3 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$, where $\rho\colon H\to(S^{1})^{h+1}$ is the symplectic slice representation at $p$. Hence, by Lemma 4.4 and the Morse Lemma, after possibly replacing $g$ by g-g, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, that $g([t,\alpha,z])=|z|^{2}$, and that $p=[1,0,0]$. Let $N$ be the degree of the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$. By Lemma 2.7, the map $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a homeomorphism; moreover, it restricts to a diffeomorphism from $({Y/\!/T})\smallsetminus\{[t,0,0]\}$ to ${\mathbb{C}}\smallsetminus\{0\}$. By Lemma 2.9, there exists a positive real number $C$ such that $g([t,\alpha,z])=C|P(z)|^{\frac{2}{N}}$ for all $t,\alpha,z]\in\Phi_{Y}^{-1}(0)$. Let $S\colon{\mathbb{C}}\to{\mathbb{C}}$ be the homeomorphism that, in polar coordinates, takes $(r,\theta)$ to $(\sqrt{C}\,r^{\frac{1}{N}},\theta)$; it restricts to a diffeomorphism ${\mathbb{C}}\smallsetminus\{0\}\to{\mathbb{C}}\smallsetminus\{0\}$. Define $\Psi\colon{Y/\!/T}\to{\mathbb{C}}\simeq{\mathbb{R}}^{2}$ by $\Psi([t,\alpha,z])=S(P(z))$. Since
@@
for all $t,\alpha,z]\in{Y/\!/T}$, the map $\Psi$ satisfies properties (1) – (3). ∎
@@ 5. The semitoric case
The goal of this section is to prove that focus-focus points satisfy the hypotheses of Proposition 3.4 (see Lemma 5.4 below). This allows us to provide a short and direct new proof of Vu Ngoc’s theorem that the fibers of semitoric systems are connected [VN] – see Theorem 5.7. In both proofs, the key arguments follow from Morse theory; the most significant difference is that we construct a new smooth structure on the reduced spaces, while he uses the natural smooth structure on $M$.
@@
Identify $T$ with $S^{1}$. By the local normal form for focus-focus points, we may assume that the integrable system is $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},\varphi\circ f_{Y})$ and $p$ is the origin, where $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},f_{Y}=(\Phi_{Y},g_{Y}))$ is as in Example 5.2 and $\varphi\colon{\mathbb{R}}\times{\mathbb{R}}\to{\mathbb{R}}\times{\mathbb{R}}$ is a diffeomorphism [C, VW]. Since by assumption the first component of $\varphi\circ f_{Y}$ induces an effective $S^{1}$-action on $Y$, and since the fibers of $f_{Y}$ are not compact, up to a translation $\Phi_{Y}$ and the first component of $\varphi\circ f_{Y}$ are equal up to sign (cf. [HSS, Proposition 3.9]). Since the integrable systems $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},f_{Y}=(\Phi_{Y},g_{Y}))$ and $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},(-\Phi_{Y},g_{Y}))$ are isomorphic, the claim follows. ∎
@@
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a 4-dimensional complexity one $T$-space. Assume that $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is focus-focus. Then there exist an open neighborhood $U$ of $p]\in{M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
@@
We are now almost ready to give our proof of Vu Ngoc’s theorem on the connectedness of the fibers of semitoric systems; however, we also need the following important result, which is proved in [At, GS2, LMTW].
@@
Let $T$ act on a connected symplectic manifold $(M,\omega)$ with moment map $\Phi\colon M\to{\mathfrak{t}}^{*}$. If $\Phi$ is proper, then every fiber of $\Phi$ is connected and $\Phi$ is open as a map to $\Phi(M)$.
@@
Let $N$ be a connected manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a proper Morse-Bott function. If no critical point of $\overline{g}$ has index or coindex one, then the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$.
@@
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be a semitoric system such that $\Phi$ is proper. Then the fiber $f^{-1}(\beta,c)=\Phi^{-1}(\beta)\cap g^{-1}(c)$ is connected for all $(\beta,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$.
@@
Fix $\beta\in{\mathfrak{t}}^{*}$. Assume without loss of generality that $\beta=0$. Since $\Phi$ is proper, the reduced space M//T{M/\!/T} is compact and connected by Theorem 5.5.
If the reduced space M//T{M/\!/T} is a single point then obviously $\Phi^{-1}(0)$ consists of a single $T$-orbit. Hence, the fiber $f^{-1}(0,c)$ is connected for all $c\in{\mathbb{R}}$.
Otherwise, since the orbit of a short point in $\Phi^{-1}(0)$ is an isolated point in $\Phi^{-1}(0)/T$, every point in $\Phi^{-1}(0)$ is tall. If $p\in\Phi^{-1}(0)$ is a critical point of $g$ modulo $\Phi$, then $p$ is also a singular point of $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$. So, by assumption, $p$ either has purely elliptic type or is focus-focus. By Propositions 3.4 and 4.5, and Lemma 5.4, M//T{M/\!/T} can be given the structure of a smooth closed, connected, oriented surface such that $\overline{g}:{M/\!/T}\to{\mathbb{R}}$ is a Morse function with no critical point of index 1. By Proposition 5.6, the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$. Since $T$ is connected, it follows that $f^{-1}(0,c)$ is also connected. ∎
@@ 6. Non-degenerate singular points
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Let $p\in\Phi^{-1}(0)$ be a tall non-degenerate singular point. By Lemma 3.2 and Proposition 4.5, if $p$ has purely elliptic type, then we can put a smooth structure on M//T{M/\!/T} near $[p]$ so that $\overline{g}$ is locally Morse, and hence $p$ satisfies the hypotheses of Proposition 3.4. The aim of this section is to extend this result to the remaining cases (see Proposition 6.3). In particular, if $p$ either has a focus-focus block, or has a hyperbolic block and disconnected $T$-stabilizer, then $[p]$ is a regular point of $\overline{g}$, while if $p$ has a hyperbolic block and connected $T$-stabilizer, then $[p]$ is a singular point with index one. We start by recalling some facts about non-degenerate singular points.
@@
Before proceeding to the tall case, we start by analyzing short points.
Lemma 6.2.
Proof.
For simplicity, assume that $g(p)=0$. By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. Let $j\colon{\mathbb{C}}^{h+1}\to T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ denote the inclusion $z\mapsto[1,0,z]$. By Lemma 3.2, $p$ is a critical point of $g$ modulo $\Phi$. Hence, the $(T\times{\mathbb{R}})$-orbit $\mathcal{O}$ through $p$ coincides with the $T$-orbit $\mathfrak{O}$ through $p$ and the Taylor polynomial $T^{1}_{0}(j^{*}g)$ of degree one at $0\in{\mathbb{C}}^{h+1}$ vanishes. Thus we may identify $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}=(T_{p}\mathfrak{O})^{\omega}/T_{p}\mathfrak{O}$ with ${\mathbb{C}}^{h+1}$ so that the homogeneous moment map for the linearized action at $p$ is the Taylor polynomial of degree two $T^{2}_{0}(j^{*}f)=(T^{2}_{0}(j^{*}\Phi),T^{2}_{0}(j^{*}g))=(\Phi_{H},T^{2}_{0}(j^{*}g))$, where $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ is the homogeneous moment map for the $H$-action.
We can now state the main result of this section.
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Assume that $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is non-degenerate and tall; let $N$ be the degree of the defining polynomial of the local model for $p$. Then $p$ has at most one non-elliptic block; furthermore,
  1. (1)
    If $p$ has a hyperbolic block and connected $T$-stabilizer, then $N=1$. Moreover, there exist an open neighborhood $U$ of $p]\in{M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
@@
      $(\overline{g}\circ\Psi^{-1})(x,y)=x^{2}-y^{2}$ for all $(x,y)\in\Psi(U)$.
@@
    If $p$ either has a focus-focus block, or has a hyperbolic block and disconnected $T$-stabilizer, then $N=2$. Moreover, there exist an open neighborhood $U$ of $p]\in{M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
    1. (a)